OpenStudy (fantageplayer):

A rectangle has sides measuring (2x + 7) units and (5x + 9) units. Part A: What is the expression that represents the area of the rectangle? Show your work to receive full credit. (4 points) Part B: What are the degree and classification of the expression obtained in Part A? (3 points) Part C: How does Part A demonstrate the closure property for polynomials? (3 points)

OpenStudy (fantageplayer):

@welshfella

OpenStudy (fantageplayer):

Part A: (2x+7)(5x+9) =(2x+7)(5x+9) =(2x)(5x)+(2x)(9)+(7)(5x)+(7)(9) =10x2+18x+35x+63 =10x2+53x+63

OpenStudy (fantageplayer):

thats what i have so far

OpenStudy (photon336):

area is length * width \[(2x+7)(5x+9)\] so we would expand this out. yep you're on the right track.

OpenStudy (welshfella):

yes

OpenStudy (fantageplayer):

thanks

OpenStudy (fantageplayer):

can you guys explain the next part

OpenStudy (fantageplayer):

im kinda confused

OpenStudy (welshfella):

the degree is the value of the highest power That is 2.

OpenStudy (fantageplayer):

oh yah i forgot

OpenStudy (fantageplayer):

thanks

OpenStudy (welshfella):

as for the classification i think in US its called a trinomial . In UK we would tend to call it a quadratic.

OpenStudy (fantageplayer):

yup

OpenStudy (fantageplayer):

i used to live in uk

OpenStudy (welshfella):

Part C - same answer as previous Q.

OpenStudy (welshfella):

OH OK! Where are you now?

OpenStudy (fantageplayer):

US

OpenStudy (fantageplayer):

used to live in england

OpenStudy (fantageplayer):

anyway i still dont get Part C

OpenStudy (welshfella):

2x + 7 and 5x + 9 are both polynomials ( in fact they are binomials) - multiplying them gives another polynomial. This demonstrates that polynomials are closed under multiplication.

OpenStudy (fantageplayer):

thanks! again

OpenStudy (welshfella):

yw